OpenStudy (anonymous):

Christa buys 14 flowerpots. Some are small and cost $2.40 each. The rest are large and cost $5.60 each. She spends a total of $49.60. Which equation models this situation? Let s represent the number of small flowerpots she buys. A. 5.6s + 2.4(s + 14) = 49.6 B. (2.4 + 9.6)(14 – s) = 49.6 C. 2.4s + 5.6(14 – s) = 49.6 D. 5.6s + 2.8(14) = 49.6

OpenStudy (redheadangel):

. you are correct. 2.let number of small small pots she buys = s total cost of small pots = 2.4s Total number of pots bought = 14 So number of large pots bought = 14-s total cost of large pots = 5.6(14-s) Total cost = 2.4s + 5.6(14-s) = 49.6 3. Let they have same amount of money in x months. Mira will have in her account = 760 - 40x Paulina will have in her account = 210 + 70x Since both have equal amount, 760 - 40x = 210 + 70x ⇒ -40x - 70x = 210 - 760 ⇒-110x = -550 ⇒ x = (-550)/ (-110) ⇒x = 5 months

OpenStudy (redheadangel):

In other words: C